# Some remarks on real algebraic maps which are topologically special generic maps

### v4

@@ Abstract.
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin. This is a diffeomorphism. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. We thus have a homeomorphism from $W_{0}$. The image of the homeomorphism is the union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=\partial{\bar{N}}_{0}\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
@@
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$.
We can define another non-singular real algebraic manifold
$M_{0,t}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t)=0\}$
for $0<t\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. This is diffeomorphic to $M_{0}$. This new set has properties similar to important ones of $M_{0}$.
@@
We go back to the trivial bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, this smooth bundle which may not be trivial in general is also trivial. Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
We investigate the bundle in Main Theorem 3 (1) and (2).
We can present a simple and natural proof. However, we present a complicated proof to respect general cases. Let the bundle in Main Theorem 3 (1) and (2) denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$. First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1 in a revised style. More precisely, Proposition 2.1 of [43] is a corresponding result. It also refers to [36] there. We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$. From $W_{0}$, we remove the total space of this bundle over ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})$. We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$. This fiber is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$. This is also thanks to Ehresmann’s fibration theorem (or a relative version of this). The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$. The structure group of the bundle $M_{{\rm B},0}$ is naturally regarded as the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$. For each $0<t\leq 1$, it is a subbundle of the following bundle: the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly. For each two obtained smooth bundles for $0<t\leq 1$, fibers are regarded as a manifold homeomorphic to $\partial D^{m-n+1}$ and a manifold homeomorphic to $D^{m-n+1}$, respectively. The former bundle is also obtained by restricting the fiber to the boundary homeomorphic $\partial D^{m-n+1}$ of course. The manifold homeomorphic to $D^{m-n+1}$ is diffeomorphic to $D^{m-n+1}$ if $m-n+1\neq 4,5$ or the boundary is diffeomorphic to $\partial D^{m-n+1}=S^{m-n}$ with the constraint $m-n+1\neq 4$ of course. We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before. This element can be regarded as an element canonically extended to an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before.
By these arguments, we may discuss the bundles and the structure groups in the following simple situation essentially. We can regard the structure group of the bundle $W_{1,t}$ as a group acting on the unit disk $D^{m-n+1}$ preserving the distance of each point and the origin $0\in D^{m-n+1}$. We explain about an element ${\psi}_{t}$ of the group. We can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}-\{0\}$ and $0<s\leq 1$. Of course 00 here is the origin and $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors. Here we may regard the relation $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$ by considering the distance $||x||$ for $x\in D^{m-n+1}-\{0\}\subset{\mathbb{R}}^{m-n+1}$.
We consider the limit at $s=0$. For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$.
These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1). We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
@@
By our arguments, we can also check Main Theorem 3 (3) easily. A result close to this is also presented in Theorem 1 (3) of [17]. ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$ there. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there.
@@
Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$. For this we can choose $a$ as an arbitrary positive real number. The bundles in (1) and (2) are regarded as ”linear bundles defined before”. Furthermore, in this case, we can consider a suitable situation and regard that ${\Phi}_{F_{S^{m-n}},a}$ is replaced by a diffeomorphism, that the homeomorphism mapping each sphere is regarded as a diffeomorphism, and that (3) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3. This is also for Main Theorem 2.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of unit spheres, studied in [11], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
@@
Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a Morse function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$. We have a special generic map $f_{0}$ as in the case of Main Theorem 2.
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with $k_{j,1}$ being positive numbers and $k_{j,2}$ being positive integers. See also Remark 3 of [14].
@@
In [14, 15, 18], we consider cases where polynomials are given beforehand. In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark. This does not affect main ingredients of [14, 15, 18]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
Related to Remark 3, we add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [25, 26, 27, 28, 29, 30, 31, 32]. See also [41]. This is also closely related to [4]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory and this mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not explain about elementary notions on graphs precisely. We except related knowledge. We also note that we omit several rigorous discussions in Remark 4 and that for our main ingredients of the present paper we do not need to understand this rigorously.
@@
  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is an smooth embedding.
@@
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of connected components of all preimages of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$ under the natural equivalence relation. Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ containing $e(K)$ and its Poincaré-Reeb graph is isomorphic to and naturally regarded as $e(K)$. Moreover, we can have the situation in the following.
@@
    The restriction of ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
@@
    $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial. This is also obtained by approximation of the Weierstrass type in the class $C^{2}$.
According to arguments similar to one in ”Special Case 5” of [20], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
@@
As another method, we can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
We give a short remark on [17] last.
@@
[17] has no essential errors and we have also found several more rigorous arguments. Some of these important arguments are discussed in the present paper. We will revise the version and submit it to the refereed journal, based on comments and our arguments and idea here. Note also that the content of [17] is independent of the content of the present paper essentially.

### v5

@@ Abstract.
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
  • If $x\in{\bar{N}}_{0}-{\rm Int}\ N(\partial{\bar{N}}_{0})$, then $g(t)=1$.
This rule is in fact same as the previous one. This gives a diffeomorphism. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a homeomorphism from $W_{0}$. The image of the homeomorphism is the disjoint union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
@@
From this we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$. The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$. In other words, it is also regarded as a ”canonical projection of $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ onto ${\rm Int}\ {\bar{N}}_{0}$”.
We can define another set
$M_{0,t^{\prime}}:=\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$
for $0<t^{\prime}\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. We present several properties based on our presented arguments. We can also know $M_{0,t^{\prime}}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$. For distinct real numbers $t^{\prime}=t_{1},t_{2}$ here, the sets $M_{0,t_{1}}-((\partial\bar{N_{0}})\times\{0\})$ and $M_{0,t_{2}}-((\partial\bar{N_{0}})\times\{0\})$ are disjoint. These sets $M_{0,t^{\prime}}-((\partial\bar{N_{0}})\times\{0\})$ are $m$-dimensional smooth manifolds with no boundaries. They are seen as the graphs of smooth functions as $M_{0}-((\partial\bar{N_{0}})\times\{0\})$ and diffeomorphic to $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. Here some variable $y_{j}$ is regarded as the variable of the target of the function. We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold. This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$.
From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}})-((\partial{\bar{N}}_{0})\times\{0\}))$.
We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). For this, first, remember that the projection is the restriction of the canonical projection ${\pi}_{m+1,n}$. Remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function. Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example. The relation $S(f_{0})=(\partial{\bar{N}}_{0})\times\{0\}$ also holds.
@@
In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle. Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
Last we consider a small isotopy as in the proof of Main Theorems 1 and 2. This completes the proof of the statement of Main Theorem 1.
We can also check the property (5) by investigating arguments of STEP 3 of our proof of Main Theorems 1 and 2 and ”Discussion 14” of [21] again. In other words, we can check that around the set $S(f_{0})$, $f_{0}$ is represented as the product map of a natural function and the identuty map on $S(f_{0})$ as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
@@
We can also check Main Theorem 3 (6) easily thanks to our arguments. A result close to this is also presented in Theorem 1 (3) of [18]. There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there.
@@
Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$. For this we can choose $a$ as an arbitrary positive real number. The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle: ${\Delta}^{k}$ is $D^{k}$ in this case. Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3. This is also for Main Theorem 2. Main Theorem 2 gives one of most important specific cases.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example. We also apply related arguments implicitly in some of our previous arguments in the present paper.
@@
We give additional explicit cases as Examples. They are also a kind of remarks.
@@
Consider the case of the form $F_{S^{m-n}}(y,t)=t-F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y)$ with a function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ satisfying the following.
  • The function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is a Morse function.
  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on a smooth closed manifold. The manifold is a non-singular real algebraic manifold of dimension $m-n$.
  • The minimum of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is 00 and the condition ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$ holds.
Example 1 is generalized here.
We have a special generic map $f_{0}$ as in the case of Main Theorems 1 and 2.
Example 3.
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example. See also Remark 3 of [15]. Some examples of Example 2 are regarded as specific cases of the present cases.
@@
In [15, 16, 18], we consider cases where polynomials are given beforehand. In short, only STEP 2 in our proof of Main Theorems 1 and 2 is considered and related short exposition on STEP 3 in our proof of Main Theorems 1 and 2 is presented where the latter presentation is a kind of short remark. This does not affect main ingredients of [15, 16, 18]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
Remark 3 is a remark on STEP 1 in our proof of Main Theorems 1 and 2. This is also independent of our present study. Here we also add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33]. See also [42]. Remark 3 is also closely related to [4]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory. This mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not present elementary notions on graphs precisely. We except related knowledge. We also note that we omit several rigorous discussions in Remark 3 and that for our main ingredients of the present paper we do not need to understand this rigorously. However, we give comments that this is also related to Remark 1 and 2.
@@
  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is a smooth embedding.
  • At distinct vertices, the values of ${\pi}_{2,1}\circ e$ are always distinct.
@@
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation is considered. Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
@@
    $e(K)\subset R_{K}$.
@@
    The Poincaré-Reeb graph of $R_{K}$ is isomorphic to and naturally regarded as $e(K)$.
  3. (3)
    We also have the situation in the following. Of course $\overline{R_{K}}$ is the closure of $R_{K}$ in ${\mathbb{R}}^{2}$.
    1. (a)
      The restriction of ${\pi}_{2,1}$ or ${\pi}_{2,1}{\mid}_{\overline{R_{K}}}:\overline{R_{K}}\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
    2. (b)
      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
A desired region $R_{K}$ can be also obtained by an approximation of the Weierstrass type in the class $C^{2}$. According to arguments similar to ones in ”Special Case 5” of [21], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
@@
We can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
The following is closely related to one of central problems in real algebraic geometry. This may be closely related to our future work.
@@
Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces has been difficult and important in real algebraic geometry. This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools. At present, even explicit construction and classifications for cases of real algebraic curves are difficult. Our pioneering studies of the construction respect unit spheres and canonical projections of them where we do not restrict attention to $1$-dimensional cases. Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.
The author recognizes that our construction shows explicit cases of Viro’s methods. In short, Viro’s methods glue local hypersurfaces together. We can have non-singular real algebraic hypersurfaces by deforming and smoothing (Viro’s patchworking). In our paper we do not need to understand this theory well. However, to study more, we need to understand the theory more.
We close our paper by a short remark on [18].
Remark 6.
[18] has no essential errors and we have also found that we need several more rigorous exposition. Some of these important arguments are discussed in the present paper. We will revise the version and submit our revised version to the refereed journal, based on the presented comments by editors and referees on [18] and our new arguments and idea in the present paper. Note also that the content of [18] is independent of the content of the present paper essentially. We have also found we can revise [18] in a self-contained way although we need to respect some of our paper. We will do so as we can.